OpenStudy (anonymous):

f[g(x)] = g[f(x)] for any two functions. (hint: Use f(x) = x + 1 and g(x) = x2). Select one: True False

OpenStudy (campbell_st):

well \[f(g(x)) = f(x^2)\] so in f(x) you replace x with x^2 so \[f(g(x)) = x^2 + 1\] now for g(f(x)) = g(x + 1) so you replace x by x + 1 \[g(f(x)) = (x + 1)^2 = x^2 + 2x + 1\] so you decide if they are the same...

OpenStudy (anonymous):

true